During basketball practice, Robert makes 5 free-throw attempts. If the probability that he gets a basket on any one free throw is 60%, independent of the other attempts, what is the probability that he makes at least 1 basket out of 5 attempts?
Answer #1

p = 0.6

q = 1 - 0.6 = 0.4

P(at least 1 basket out of 5) = 1 - P(no basket) = 1 - 0.4^5 = 0.98976

Hence, 0.98976 is the probability that he makes at least 1 basket out of 5 attempt.

If it ask for 4 decimal places then answer is 0.9898

Please comment if any doubt. Thank you.